. The hospital record shows that the mean weight of newly born baby is greater than 7 lbs. A researcher takes a sample of 25 newly born babies and found to have a mean weight of 6.97lbs with the standard deviation of 0.77 lbs. Test the claim at 0.05 level of significance.




1
Expert's answer
2022-05-26T08:13:32-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 7"

"H_1:\mu>7"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=24" and the critical value for a right-tailed test is "t_c =1.710882."

The rejection region for this right-tailed test is "R = \{t:t>1.710882\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{6.97-7}{0.77/\sqrt{25}}=-0.1948"

Since it is observed that "t=-0.1948<1.710882=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for right-tailed, "df=24" degrees of freedom, "t=-0.1948" is "p=0.576405," and since "p=0.576405>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is greater than 7, at the "\alpha = 0.05" significance level.